Find all real or imaginary solutions to the following equation, b^4+13b^2+36=0
Answer by stanbon(75887)

b^4+13b^2+36=0
------------------
Factor to get:
(b^2+4)(b^2+9) = 0
b^2=-4 or b^2=-9
b = 2i or b = -2i
b = 3i or b = -3i
================
Cheers,
Stan H.